### v1

@@ 1 Introduction
Represented by ChatGPT [^1] https://chatgpt.com, the rise of large language models (LLMs) has showcased impressive language capabilities and professional competence, as well as providing correct, polite, and knowledgeable responses, which is surprising and admirable. This performance should be largely attributed to the preference alignment process, a currently necessary step that LLMs undergo before public deployment, to prevent LLMs from potentially generating offensive, toxic, or misleading content.
@@
CriticGPT (mcaleese2024llmcriticshelpcatch), Generative Verifier (zhang2024generativeverifiersrewardmodeling) , leaf, text width=22.0em ] ] ] ] [ Optimization (§6) [ Pointwise method [ RFT (rft; star; raft), PPO (schulman2017proximal), GRPO (shao2024deepseekmath), ReMax (li2023remax), KTO (ethayarajh2024kto) , leaf, text width=33.6em ] ] [ Pairwise contrast [ CoH (liu2024chain), SLiC (zhao2022calibrating), DPO (rafailov2024direct), IPO (azar2023general), Sr-DPO (yu2024direct), ORPO (hong2024orpo),
Mallow-DPO (chen2024mallows), GRPO (ramesh2024group), DPO-positive (pal2024smaug), CPL (hejna2023contrastive), EXO (ji2024towards),
@@
VCB (mao2024don), LiPO (liu2024lipo), LIRE (zhu2024lire) , leaf, text width=33.6em ] ] [ Training-Free [ Input Optimization [ BPO (cheng2023black), URIAL (lin2023unlocking), OPO (xu2023align) , leaf, text width=22.0em ] ] [ Output Optimization [ ICDPO (song2024icdpo), Aligner (ji2024aligner), RAIN (li2023rain),
@@
This new framework pivots on two key insights: First, the distinction between on-policy and off-policy settings essentially depends on different sources of data, which can be decoupled from algorithms like PPO or DPO. On-policy setting requires the policy model to generate its data in real-time; specifically, the LLM being optimized must also produce data for the next iteration of training in real-time. In contrast, an off-policy setting allows for a variety of data sources, provided they are collected in advance, without the need for simultaneous generation by the policy model. Many current works employ the transition of specific algorithms between on-policy and off-policy settings (guo2024direct; shao2024deepseekmath). Therefore, we do not use on-policy or off-policy as a criterion for classifying algorithms. Second, inspired by existing work (shao2024deepseekmath), the essence of the objective of optimization in reinforcement learning and supervised finetuning-based methods are actually quite similar. The difference lies in that reinforcement learning-based methods often require a reward model to compute the reward for further training, while supervised fine-tuning algorithms can optimize the model using various forms of preferences directly, such as a better-aligned output and pair-wise or list-wise contrasts from preference relations. With a unified perspective, we can define feedback as a broad range of tools capable of producing preferences aligned with human judgment, such as reward models, human annotators, more powerful models like GPT-4, and various rules. Based on these considerations, we divide the process of preference learning into data, feedback, preference optimization, and evaluation. The taxonomy of our paper is shown in Figure 2.
In summary, our paper investigates and organizes existing work related to preference learning for LLM, offering a unified and novel perspective. Further, based on the content of this review, we summarize several future directions in this area, intending to bring insights for further research.
@@ 3       $\pi_{ref}\leftarrow\pi_{\theta}$;
9             $\mathcal{B}\leftarrow$ Sample response from $\pi_{\theta}$ using $\mathcal{Q}$ and get the preference siganl in real time;
@@
12             $\mathcal{B}\leftarrow$ Get a batch of data with preference siganl from the pre-stored $\mathcal{D}$;
13
@@ 4.1 On-policy Data Collection
For problems that involve multi-step solutions, there has also been some research (wang2024mathshepherd; zhang2024accessinggpt4levelmathematical; zhang2024restmctsllmselftrainingprocess; feng2024alphazeroliketreesearchguidelarge; luo2024improvemathematicalreasoninglanguage; qi2024mutualreasoningmakessmaller; zhu2024recoveringmentalrepresentationslarge) employing Monte Carlo tree search (MCTS) KocsisMCTS to enhance the diversity and performance of the data sampling. MCTS originated from the advancements made in AlphaGo. The fundamental concept of MCTS involves evaluating various strategies through numerous simulations, or rollouts, to determine which strategy yields superior results. This approach resembles a methodical and deliberative thinking process, contrasting with greedy decoding methods that prioritize immediate gains. The core operations of MCTS can be categorized into four distinct phases: selection, expansion, simulation, and backpropagation. MCTS’s efficient search strategy enables models to generate higher-quality data while simultaneously acquiring step-level labels. This refined data can subsequently be utilized to enhance model performance during decoding (qi2024mutualreasoningmakessmaller; zhu2024recoveringmentalrepresentationslarge; feng2024alphazeroliketreesearchguidelarge), train reward models (wang2024mathshepherd; luo2024improvemathematicalreasoninglanguage), and fine-tune the model (zhang2024accessinggpt4levelmathematical; zhang2024restmctsllmselftrainingprocess; feng2024alphazeroliketreesearchguidelarge).
@@ 4.2 Off-policy Data Collection
Off-policy data collection entails gathering training data independently of the LLM’s learning process. This method is generally easier than on-policy data collection, largely due to the availability of open-source preference datasets. Alternatively, we can also compile a dataset in advance using the initial model $\pi_{\theta^{0}}$. The off-policy data collection strategy ensures a more diverse training dataset that can often yield improvements in the LLM’s preference learning process. There are two main sources of preference data, the ones from human annotators and those generated by more advanced LLMs.
@@
SHP pmlr-v162-ethayarajh22a collects 385K human preferences over responses to questions/instructions in 18 domains for training RLHF reward models and NLG evaluation models.
@@ 6 Algorithms
Preference learning algorithms update LLMs to align with human preferences based on data and rewards. In this section, we categorize the preference algorithms into four groups: point-wise methods, pair-wise contrasts, list-wise contrasts, and training-free alignment. For some representative algorithms belonging to pair-wise/list-wise contrasts, we also provide their detailed loss designs in Table 1.
6.1 Point-wise Methods
@@
where $\pi_{ref}$ is the supervised fine-tuned model and $\pi_{\theta}$ is initialized as $\pi_{ref}$. $\beta$ is the KL divergence coefficient, which controls the deviation from the original model. Despite the impressive performance demonstrated by OpenAI, the PPO algorithm requires extensive computational resources and suffers from sample inefficiency. Another drawback of the PPO algorithm is its training instability, making it challenging to determine the appropriate hyperparameters for the PPO method. Some alternative point-wise methods have introduced enhancements to the PPO framework. For instance, ReMax (li2023remax) and GRPO (shao2024deepseekmath) eliminate the use of the critic model that PPO relies on, thereby reducing the consumption of storage resources during training.
@@
6.2 Pair-wise Contrasts
@@
6.3 List-wise Contrasts
@@
Vanilla list-wise contrasts may lead to performance degradation due to biased estimation of candidates, requiring more fine-grained design. wang2023making and mao2024don implement multiple calibration strategies on computed scores for different ranking objectives to alleviate over-fitting. Differently, liu2024lipo and zhu2024lire opt to design re-weighting mechanisms on each term in loss functions with external scoring information, which are beneficial to precise scoring ability of $\pi$.
@@ 7 Evaluation
The methods for evaluating alignment are divided into two parts: rule-based evaluation as Sec. 7.1 and LLM-based evaluation as Sec. 7.2.

### v2

@@ 1 Introduction
Represented by ChatGPT [^1] https://chatgpt.com, the rise of large language models (LLMs) has showcased impressive language capabilities and professional competence, as well as providing correct, polite, and knowledgeable responses, which is surprising and admirable. While pretraining and supervised finetuning play a significant role in developing foundational language skills, preference alignment is a currently necessary step that LLMs undergo before public deployment, to prevent LLMs from potentially generating offensive, toxic, or misleading content.
@@
CriticGPT (mcaleese2024llmcriticshelpcatch), Generative Verifier (zhang2024generativeverifiersrewardmodeling) , leaf, text width=22.0em ] ] ] ] [ Optimization (§6) [ Pointwise method [ RFT (rft), RAFT (raft), Star (star), PPO (schulman2017proximal), ReMax (li2023remax), KTO (ethayarajh2024kto) , leaf, text width=33.6em ] ] [ Pairwise contrast [ CoH (liu2024chain), SLiC (zhao2022calibrating), DPO (rafailov2024direct), IPO (azar2023general), Sr-DPO (yu2024direct), ORPO (hong2024orpo),
Mallow-DPO (chen2024mallows), GRPO* (ramesh2024group), DPO-positive (pal2024smaug), CPL (hejna2023contrastive), EXO (ji2024towards),
@@
VCB (mao2024don), LiPO (liu2024lipo), LIRE (zhu2024lire), GRPO (shao2024deepseekmath) , leaf, text width=33.6em ] ] [ Training-Free [ Input Optimization [ BPO (cheng2023black), URIAL (lin2023unlocking), OPO (xu2023align) , leaf, text width=22.0em ] ] [ Output Optimization [ ICDPO (song2024icdpo), Aligner (ji2024aligner), RAIN (li2023rain),
@@
This new framework pivots on two key insights: First, the distinction between on-policy and off-policy settings essentially depends on different sources of data, which can be decoupled from algorithms like PPO or DPO. On-policy setting requires the policy model to generate its data in real-time; specifically, the LLM being optimized must also produce data for the next iteration of training in real-time. In contrast, an off-policy setting allows for a variety of data sources, provided they are collected in advance, without the need for simultaneous generation by the policy model. Many current works employ the transition of specific algorithms between on-policy and off-policy settings (guo2024direct; shao2024deepseekmath). Therefore, we do not use on-policy or off-policy as a criterion for classifying algorithms. Second, inspired by existing work (shao2024deepseekmath), the essence of the objective of optimization in reinforcement learning and supervised finetuning-based methods are actually quite similar. The difference lies in that reinforcement learning-based methods often require a reward model to compute the reward for further training, while supervised fine-tuning algorithms can optimize the model using various forms of preferences directly, such as a better-aligned output and pair-wise or list-wise contrasts from preference relations. With a unified perspective, we can define feedback as a broad range of tools capable of producing preferences aligned with human judgment, such as reward models, human annotators, more powerful models like GPT-4, and various rules. Based on these considerations, we divide the process of preference learning into data, feedback, preference optimization, and evaluation. The taxonomy of our paper is shown in Figure 2. Moreover, we provide clear running examples of some common algorithms within this framework to facilitate the readers’ understanding of the algorithms, which are shown in Figure 3 and Figure 4.
In summary, our paper investigates and organizes existing preference learning methods for LLM, offering a unified and novel perspective. Further, based on the content of this survey, we summarize several future directions in this area, intending to bring insights for further research.
@@ 3       $\pi_{ref}\leftarrow\pi_{\theta}$;
9             $\mathcal{B}$leftarrow Sample response from $\pi_{\theta}$ using $\mathcal{Q}$;
10             $\mathcal{R}$leftarrow Get the feedback from the environment $\mathcal{E}$ in real time;
11
@@ 10             $\mathcal{R}$leftarrow Get the feedback from the environment $\mathcal{E}$ in real time;
14             $\mathcal{B}\leftarrow$ Get a batch of data with the preference feedback from the pre-stored $\mathcal{D}$;
15
@@ 4.1 On-policy Data Collection
For problems that involve multi-step solutions, there has also been some research (wang2024mathshepherd; zhang2024accessinggpt4levelmathematical; zhang2024restmctsllmselftrainingprocess; feng2024alphazeroliketreesearchguidelarge; luo2024improvemathematicalreasoninglanguage; qi2024mutualreasoningmakessmaller; zhu2024recoveringmentalrepresentationslarge; xiong2024watchstepllmagent) employing Monte Carlo tree search (MCTS) KocsisMCTS to enhance the diversity and performance of the data sampling. MCTS originated from the advancements made in AlphaGo. The fundamental concept of MCTS involves evaluating various strategies through numerous simulations, or rollouts, to determine which strategy yields superior results. This approach resembles a methodical and deliberative thinking process, contrasting with greedy decoding methods that prioritize immediate gains. The core operations of MCTS can be categorized into four distinct phases: selection, expansion, simulation, and backpropagation. MCTS’s efficient search strategy enables models to generate higher-quality data while simultaneously acquiring step-level labels. This refined data can subsequently be utilized to enhance model performance during decoding (qi2024mutualreasoningmakessmaller; zhu2024recoveringmentalrepresentationslarge; feng2024alphazeroliketreesearchguidelarge), train reward models (wang2024mathshepherd; luo2024improvemathematicalreasoninglanguage), and fine-tune the model (zhang2024accessinggpt4levelmathematical; zhang2024restmctsllmselftrainingprocess; feng2024alphazeroliketreesearchguidelarge).
@@ 4.2 Off-policy Data Collection
Off-policy data collection entails gathering training data independently of the LLM’s learning process. This method is generally easier than on-policy data collection, largely due to the availability of open-source preference datasets. Alternatively, we can also compile a dataset in advance using the initial model $\pi_{\theta^{0}}$. The off-policy data collection strategy ensures a more diverse training dataset that can often yield improvements in the LLM’s preference learning process. There are two main sources of preference data, the ones from human annotators and those generated by more advanced LLMs. Please note that as relevant research continues to advance, the number of open-source datasets related to preference learning is increasing. Consequently, it is challenging to compile a comprehensive list of all datasets. Therefore, we will only highlight a few representative works.
@@
SHP pmlr-v162-ethayarajh22a consists of 385K human preferences regarding responses to questions in 18 subject areas, reflecting user preferences for helpfulness. In contrast to HH-RLHF, SHP relies solely on human-written data, allowing for complementary distributions between the two datasets.
@@ 6 Algorithms
Preference learning algorithms optimize LLMs to align with human preferences based on data and feedback. Based on the number of samples needed to compute the gradient coefficient in formula 1, we can categorize these algorithms into three types: point-wise methods, pair-wise contrasts, and list-wise contrasts. Point-wise methods rely on the quality of a single sample to determine the gradient coefficient, pair-wise contrasts require comparisons between pairs of samples, and list-wise contrasts involve evaluating entire lists of samples to compute the gradient coefficient. In addition to this, there are also algorithms that optimize the models without the need for training, we categorize them into training-free alignment.
In summary, we categorize the preference algorithms into four groups: point-wise methods, pair-wise contrasts, list-wise contrasts, and training-free alignment. For some representative algorithms belonging to pair-wise/list-wise contrasts, we also provide their detailed loss designs in Table 1.
6.1 Point-wise Method
@@
where $\pi_{ref}$ is the supervised fine-tuned model and $\pi_{\theta}$ is initialized as $\pi_{ref}$. $\beta$ is the KL divergence coefficient, which controls the deviation from the original model. Despite the impressive performance demonstrated by OpenAI, the PPO algorithm requires extensive computational resources and suffers from sample inefficiency. Another drawback of the PPO algorithm is its training instability, making it challenging to determine the appropriate hyperparameters for the PPO method.
To address the drawbacks of PPO outlined earlier, several alternative point-wise methods have been proposed. One such method is ReMax (li2023remax), which draws on concepts from the REINFORCE algorithm (williams1992simple). Notably, ReMax modify the calculation of gradient coefficient by incorporating a subtractive baseline value, and the baseline value can be defined as the reward of a greedy sampling response. The authors suggest that this approach reduces computational requirements by eliminating the need for a critic model, which is essential for PPO, while also promoting more stable training. From our perspective, ReMax should be "pairwise" as it introduces an additional subtractive baseline for computing the gradient coefficient. We introduce ReMax in this section to help readers understand the process without being too abrupt.
@@
6.2 Pair-wise Contrast
@@
6.3 List-wise Contrast
@@
Vanilla list-wise contrasts may lead to performance degradation due to biased estimation of candidates, requiring more fine-grained design. wang2023making and mao2024don implement multiple calibration strategies on computed scores for different ranking objectives to alleviate over-fitting. Differently, liu2024lipo and zhu2024lire opt to design re-weighting mechanisms on each term in loss functions with external scoring information, which are beneficial to precise scoring ability of $\pi$. Some alternative list-wise methods have introduced enhancements to the currently most successful PPO algorithm. For instance, GRPO (shao2024deepseekmath) samples a list of responses for a given query and employs a reward model to evaluate each output. The reward for each response is normalized by subtracting the average reward of the response list and dividing it by the list’s standard deviation. For each output, GRPO sets the advantage to the normalized reward, eliminating the use of the critic model that PPO relies on, thereby reducing the consumption of storage resources during training.
@@ 7 Evaluation
For evaluation of preference learning, the ideal approach is human assessment, such as the Chatbot Arena chiang2024chatbotarenaopenplatform, a benchmarking platform for large language models (LLMs) that facilitates anonymous, randomized matchups through crowd-sourcing. However, due to resource constraints and the potential biases associated with human evaluations, the prevailing method remains automated assessment, which is divided into two parts: rule-based evaluation as Sec. 7.1 and LLM-based evaluation as Sec. 7.2.
@@ 9 Conclusion
Appendix B The discussion about ReMAX and GRPO
In this chapter, we provide a detailed clarification of our discussion regarding the classification of the ReMAX (li2023remax) and GRPO (shao2024deepseekmath) algorithms. Both of these approaches aim to forego the critic model of PPO. From the perspective of the gradient coefficient, ReMax requires an additional reward from a sequence decoded greedily to serve as a baseline, while GRPO estimates the baseline from group scores. This makes ReMax a pairwise method and GRPO a listwise method.
However, from the standpoint of loss calculation, once both algorithms compute the reward or advantage, they utilize samples to optimize the model in a pointwise manner. This is because, unlike algorithms like DPO that use two or a group of samples to compute the loss itself, they focus solely on calculating the gradient coefficient. For the sake of fluency and readability in the main text, we did not include these considerations and simply categorized ReMAX as pointwise and GRPO as listwise.